OpenStudy (anonymous):

The sum of a number and 6 times its reciprocal is 7. Find the number. If the problem has two solutions, write your answers in the box below.

OpenStudy (anonymous):

(x + 6)*(1/x) = 7 move the 1/x to the other side by multiplying both sides *x x+6 = 7*x move all of the 'x' factors to the same side by subtracting x from both sides. 6 = 6*x divide both sides by 6 to find your answer. x = 1